Use properties of exponents to simplify. Write with positive exponents.
x^3/7 * x^1/5 / x^-1/3
I am thinking we multiply the num/ dem. But I am not sure where to start honestly. Thanks!

Answer by oscargut(686)

x^3/7 * x^1/5 / x^-1/3 = x^(3/7+1/5)/x^-1/3
x^(15/35+7/35)/x^-1/3=x^(22/35)/x^-1/3=x^(22/35 + 1/3)=x^(66/105 + 35/105)=
x^(101/105)